OpenStudy (anonymous):

x^2+14x+38 / x+8????? Please help!!!

Nnesha (nnesha):

factor quadratic equation x^2 + 14 x+38 find two number if you multiply them you should get 38 andif you add or subtract them you should get 14

Nnesha (nnesha):

btw that's not factorable so you have to use quadratic formula hmmm

OpenStudy (solomonzelman):

it is not factor-able, since the discriminant is not a perfect square. \(\large\color{slate}{ (14)^2-(4)(1)(38) }\) \(\large\color{slate}{ 44 }\)

OpenStudy (anonymous):

undefined the answer is x^2+14x +38/x+8

OpenStudy (solomonzelman):

But you can make it, \(\large\color{black}{ \displaystyle \frac{ x^2+14x+38 }{x+8} }\) \(\large\color{black}{ \displaystyle \frac{ x^2+14x+48-10 }{x+8} }\) \(\large\color{black}{ \displaystyle \frac{ x^2+14x+48 }{x+8} -\frac{ 10 }{x+8} }\)

OpenStudy (solomonzelman):

that is for polynomial division, if you don't like setting it in long or synthetic division.

OpenStudy (solomonzelman):

Or, do you want to solve for x? then shouldn't it be equal to something?

Nnesha (nnesha):

divide..........long division or synthetic divison

OpenStudy (solomonzelman):

or the way I did

OpenStudy (solomonzelman):

(I didn't finish it of course)

OpenStudy (solomonzelman):

for me?

OpenStudy (solomonzelman):

lol

Nnesha (nnesha):

Nope.